Grade 11Algebra

Solve.I x-1l^[(lnlxl)^2 -lnx^2]=|x-1l^3. And next question :-) √log (5x-x^2)/4 with base 1/4.find x

Sigma Samantasinghar
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
 
Again i am answering,
 
Case 1:
when x
U can equate Lnlxl)^2-Lnx^2 =3
or let logx = t
t^2 – 2t – 3 = 0
it has roots 
t = -1 and t = 3
thus,
logx = -1 and logx = 3
and
x = 1/e  
for x = e^3 solution rejected as e^3 is > 1.
 
Case 2: when x> 1
Again, t will have two roots as:
t = -1 and t = 3
x = 1/e  and x = e^3
For x > 1 1/e is rejected and x = e^3 is accepted.
 
Now intersection for both cases dnot satisfies them simulatneously is something looks not utterabble.
But u can observe logically,
By intersection i mean to say:
one solution for one case is being get rejected for other case.
and same for vice versa.
Thus  (Lnlxl^2)-Lnx^2 =3
cannot be equated.


Plz. post your next question as another post again!